OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-4, -3) and (-6, 7).

OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-4, -3) and (-6, 7).

OpenStudy (anonymous):

use m = (y2-y1) / (x2-x1)to get your slope. then plug your slope in for m of y=mx+b and solve for the y intercept

OpenStudy (anonymous):

roxygurl i really dont know how to do this to be truthful with you I need the answer

OpenStudy (anonymous):

haha it's alright, these kinda get confusing. it's y=-5x-23

OpenStudy (anonymous):

they want me too show all the work as in how did i arrive to this equation

OpenStudy (mimi_x3):

m = 7+3/-6+4 = -5 y = mx +b -3= -5(-4)+ b b = -23 y = -5x -23

OpenStudy (anonymous):

thank you so much mimi i just dont get this at all and if I fail this test i get kicked out of college

OpenStudy (mimi_x3):

yw (: Your in college and you don't know how to do this ?

OpenStudy (anonymous):

Do you want me to try to explain it to you again? It's really crucial that you need to know how to do something like this.

OpenStudy (anonymous):

its been a while had a bad car accident and lost like half of my memory

OpenStudy (anonymous):

yes

OpenStudy (mimi_x3):

Wasn't my explanation thorough enough ?>

OpenStudy (anonymous):

okay, so slope intercept is y=mx+b, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so when you're given two points, you have to first find the slope of the line to complete your equation. you use m= (y2-y1) / (x2-x1)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

All this is saying is you take the y in the first point and subtract it from the y in the second point and divide it by the first x point by the second x point. so, you're using both equations to find the slope

OpenStudy (anonymous):

here y I chossen the name just call me stupid i cant remember long enough

OpenStudy (anonymous):

haha it's alright, math is hard!

OpenStudy (anonymous):

so after you have your slope, plug it into your y=mx+b equation. then you'd have y=5x +b, because m represents your slope.

OpenStudy (anonymous):

now, you need to solve for the y intercept, which is represented by the y.

OpenStudy (anonymous):

its only math that i have the trouble with the numbers get screwed up in my head for some unknown reason and the kicker is i am doing computers aint that a blip.

OpenStudy (anonymous):

to do this, you'll need to use one of your coordinates. it doesn't matter which one.let's use (-6, 7). You'd plug in -6 for x and 7 for y. So, your equation becomes 7=5(-6) + y

OpenStudy (anonymous):

thank you so much your a god send

OpenStudy (anonymous):

SInce you're solving for y, you'd just multipy -6 and 5, then subtract it from both sides for it to cancel out and get y alone. You're left with y= -23. Then you just plug that into your equation, so you'll have y=5x-23

OpenStudy (anonymous):

no problem! if you need any more help, just ask!

OpenStudy (anonymous):

oh thank you sweety i think i have to more question lucky god sent me you huh

OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-5, -8) and is perpendicular to 10x – 6y = -11

OpenStudy (anonymous):

haha. well my boyfriend's a math major so he kinda taught me everything!

OpenStudy (anonymous):

wow i need to find me a girlfriend like that then? lucky you.

OpenStudy (anonymous):

haha!